FOURIER TRANSFORM MILLIMETER-WAVE SPECTROSCOPY OF THE ETHYL RADICAL IN THE ELECTRONIC GROUND STATE

\underline{EUNSOOK KIM}, HIDETA HABARA, and SATOSHI YAMAMOTO, Department of Physics and Research Center for the Early Universe, The University of Tokyo, Bunkyo-ku, Tokyo 113-0033, Japan.

The 1_{01}-0_{00} rotational transition of the ethyl radical has been detected for the first time with the Fourier transform millimeter-wave (FTMW) spectrometer. The ethyl radical is produced by discharging the C_{2}H_{5}I gas diluted in Ar. We have observed 28 paramagnetic lines in the frequency region from 43680 MHz to 43780 MHz. These lines were observed by using C_{2}H_{5}Br, CH_{3}CH_{2}COCH_{2}CH_{3}, and C_{2}H_{6} instead of C_{2}H_{5}I, although the line intensities were weaker than the C_{2}H_{5}I case. From this evidence, we concluded that the spectral lines observed are those of C_{2}H_{5}. The lines show a very complicated pattern of the fine and hyperfine structures of the doublet radical with the nuclear spins of five protons. Therefore measurements of the Zeeman effect were useful to assign the fine and hyperfine components. As a result, the most lines were ascribed to the transitions in the A2^{′′} level. The rotational constant, the spin-rotation constant, and hyperfine constants are determined by the least-squares fit.